### v2

# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

Abstract
We introduce EMSYNC, a video-based symbolic music generation model that aligns music with a video’s emotional content and temporal boundaries. It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues. We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate upcoming video scene cuts and align generated musical chords with them. Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances. We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs. Our method outperforms state-of-the-art models in objective and subjective evaluations across different video datasets, demonstrating its effectiveness in generating music aligned to video both emotionally and temporally.
1 Introduction
The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software (Falkowski-Gilski and Uhl 2020). A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement (Buhler 2018). However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube. Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content. Automatic video-based music generation offers a promising solution to this problem (Di et al. 2021; Zhuo et al. 2023; Kang, Poria, and Herremans 2024).
Existing approaches that generate music as audio waveforms lack editability (Lam et al. 2024; Copet et al. 2024). These methods compress all stages of music production, including composition, performance, editing, mixing, and mastering, into a single process, which limits creative control for musicians and producers. By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility. MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities. Musicians and producers can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs). We use “MIDI” and “music” interchangeably, as MIDI is our exclusive format. A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets. Existing datasets are either domain-specific, such as pianist hand videos (Koepke et al. 2020), or contain limited samples, around 1k (Di et al. 2021; Zhuo et al. 2023). To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset (Raffel 2016), the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs. Drawing from musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment (Deutsch 2007) and temporal synchronization (Prendergast 1992). We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries. While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model (Russell 1980). Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music. We also use a pretrained video emotion classifier to estimate discrete emotion probabilities, map them to valence-arousal values, and condition our music generator accordingly.
Temporal conditioning in MIDI-generating models presents unique challenges. While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate (Arnab et al. 2021; Liu et al. 2022). In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated (Oore et al. 2020). In event-based MIDI encoding, two primary token types exist: “note” and “time shift”. Note tokens represent pitch, while time shift tokens advance the time axis, capturing both note durations and silences. Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time. The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship. Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids (Di et al. 2021; Kang, Poria, and Herremans 2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
The selection of temporal features for video-music synchronization is also non-trivial. Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density (Di et al. 2021). Our inspection of the output samples from this approach identified two key limitations. First, the continuously varying note density can disrupt rhythmic consistency, making the underlying beat unstable and harder to follow. Second, because the mapping is dense, synchronization between music and video is difficult to perceive unless there are strong contrasts in motion speed and note density across consecutive sections. Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by sparse temporal boundaries of scene cuts (Cohen 1990; Prendergast 1992). To enable our music generator to anticipate future boundaries and align its output accordingly, we propose using boundary offsets. We conduct objective and subjective evaluations to measure the temporal synchronization between input video and generated music, and demonstrate that our approach outperforms dense temporal conditioning.
In addition, we tackle the challenge of using emotions as an intermediary to link music generation to video. Emotions are commonly represented with either categorical (Ekman 1971) or dimensional (continuous-valued) (Russell 1980) models. Based on prior findings from psychology and emotion research (Russell and Mehrabian 1977), we provide a mapping between categorical and dimensional emotion representations. This mapping enables the combined use of different emotion representations across data modalities, such as categorical emotions for video classification and dimensional emotions for MIDI generation, as in our case. This approach ensures flexibility and supports the use of training data labeled with either type of emotion representation.
We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets, where it outperforms the state-of-the-art methods. Our main contributions are as follows:
  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
  • We introduce boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
  • Based on psychological research, we establish a mapping between discrete emotion categories and continuous valence-arousal values, an enabling emotional connection between the input video and the generated music.
2 Related work
While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances. The Foley Music model (Gan et al. 2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network (Kipf and Welling 2017) and a Transformer (Vaswani et al. 2017). Similarly, Koepke et al. (2020) and Su et al. (2020) use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards. Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples (Zhao et al. 2018; Li et al. 2018; Koepke et al. 2020; Su, Liu, and Shlizerman 2020). The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music (Su, Liu, and Shlizerman 2021).
Some studies explore the more general task of generating symbolic music for arbitrary videos. The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing (Di et al. 2021). CMT employs the Lakh Pianoroll Dataset (Dong et al. 2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration (Hsiao et al. 2021). While this representation reduces sequence length, it significantly increases input dimensionality. Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute. Furthermore, note events align precisely with beat subdivisions, whereas human performance naturally deviates from beat subdivisions due to expressive timing (Fabian, Timmers, and Schubert 2014). In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8 ms resolution. This resolution is sufficient to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions (Oore et al. 2020). Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
In a follow-up to CMT, Zhuo et al. introduced the Symbolic Music Videos dataset, containing video-MIDI pairs (2023). They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples (Hawthorne et al. 2018). Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features (Radford et al. 2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application (Kang, Poria, and Herremans 2024). The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key. Their encoder-decoder transformer takes low- and high-level video features, along with user-provided primer chord and key, to generate chord sequences. These sequences are then arpeggiated using fixed patterns to create the final MIDI output. However, this method’s reliance on a fixed time grid omits the subtle expressive timing found in human performances. Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
3 Methodology
In this section, we present our video-based symbolic music generator. We match the output music to the input video based on emotions and temporal boundaries. Since there are no datasets where videos are paired with symbolic music, we employ a two-stage approach, i.e., use independent modules for video analysis and conditional music generation, and then combine them. Our pipeline is shown in Figure 1. We first highlight our music generator and how it is conditioned on emotions and temporal boundaries. We then describe how we incorporate temporal and emotional video features into our conditional music generator.
3.1 Conditional music generator
We extend an existing emotion-based MIDI generator trained with valence-arousal labels (Sulun, Davies, and Viana 2022) assigned to the Lakh MIDI dataset (Raffel 2016). By introducing temporal boundary conditioning and emotion mapping, we align music generation with input videos at inference time. Figure 2 illustrates the music generator, with the emotion-conditioning mechanism shown in the upper part and the boundary-conditioning mechanism in the lower part.
Figure 1: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values.
Training dataset and preprocessing
We train our music generator on the Lakh Pianoroll Dataset (LPD) (Dong et al. 2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset (Raffel 2016). We tokenize the pianorolls using an event-based symbolic music representation (Oore et al. 2020). Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end. These tokens also encode pitch and instrument information. For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60. Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings (Dong et al. 2018). However, our method is adaptable to datasets with different instrument groupings.
TIMESHIFT tokens are used to move along the time axis, representing both note durations and the silences between them. Each token specifies a time increment in milliseconds. For example, an 800 ms shift is encoded as TIMESHIFT_800. We use a temporal resolution of 8 ms with a maximum shift of 1000 ms. Longer durations are represented using multiple consecutive TIMESHIFT tokens. We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches. We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions (Bharucha and Krumhansl 1983). However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
To fully exploit available training data, we implement several strategies. Since we aim to generate multi-instrument compositions, we prioritize pieces with three or more instruments. However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset (Raffel 2016). As these samples still contain valuable information on musical structure, we avoid filtering them out. Instead, we prepend special tokens: FEWER_INSTRUMENTS for songs with two or fewer instruments and MORE_INSTRUMENTS for those with three or more. These tokens allow users to specify instrumentation preferences at inference time while leveraging the entire dataset during training. Additionally, to utilize MIDI samples without valence-arousal labels, we allow the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation. If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
Figure 2: Our music generator. Numbers underneath valence and arousal are sample values.
Temporal boundary matching
Our model generates music that is synchronized with the video’s temporal boundaries. We define temporal boundaries as scene cut locations in video and long-duration chord locations in music. Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries. To achieve this, we label the chords in the Lakh Pianoroll Dataset. During training, we consider only guitar and piano chords with at least three simultaneous notes that last for a minimum of two beats. These chords are labeled by inserting a CHORD token before the first ON token of each chord. To allow the model to generate chords not only at video scene cut locations but also independently where musically appropriate, we randomly remove 20% of CHORD tokens during training. An example of a labeled chord is shown in the top row of Figure 3. In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following (Bharucha and Krumhansl 1983). During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt. This occurs because the model, having no prior knowledge of the upcoming chord, may generate preceding notes that do not align naturally with it. To address this, we develop a method that enables the model to “anticipate” upcoming chords and generate preceding notes and time shifts accordingly. Additionally, we train the model to generate the CHORD token itself, ensuring rhythmic consistency in the generated music. To achieve this, we define boundary offsets for each input token, representing the remaining time until the next boundary. These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones. Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music. Sample offsets are illustrated in the bottom row of Figure 3.
We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings (Wang et al. 2020) along the feature dimension, as shown in the lower part of Figure 2. This architectural choice is motivated by several key factors. First, similar to the emotion-conditioning mechanism, we inject boundary offset encodings at the input level rather than within the transformer body, ensuring that the core model remains unchanged. This allows the model to process inputs with or without boundary offsets, enabling seamless fine-tuning by repeating the learned positional encodings along the feature dimension when boundary offsets are absent. Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two. Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them (Haviv et al. 2022; Kazemnejad et al. 2024). Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings. The resulting vector sequence consists of positional encodings augmented with boundary offset encodings. Following the standard transformer model, we add this sequence to the token embeddings (Vaswani et al. 2017) before passing it to the transformer body with relative global attention (Huang et al. 2018). The transformer’s input is represented as:
$\begin{split}\mathbf{X}_{\text{trans}}&=\text{Concat}_{s}\left(\text{FFN}_{v}(x_{v}),\text{FFN}_{a}(x_{a}),\text{Embedding}(\mathbf{x}_{t})\right)\\ &+\text{Concat}_{f}(\text{FFN}_{b}(\mathbf{b})+\mathbf{W_{pe}})\end{split}$
where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the input boundary offsets, and $\mathbf{W_{pe}}$ is the learned positional encoding. The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively. The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively.
Figure 3: Graphical illustration of boundaries. Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training. Middle: In video, scene cuts serve as video boundaries and are used during inference. Bottom: Boundary offset represents the temporal distance to the next boundary. These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation. The model generates music autoregressively, producing one token per forward pass. We maintain a time cursor by tracking the generated TIMESHIFT tokens and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token. If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary. The boundary with a difference smaller than the predefined sensitivity threshold of 1 s is considered successfully generated. We then remove these boundaries from future offset calculations by replacing them with infinity. Next, we compute the boundary offset for the generated token, regardless of its type. This offset is determined as the distance to the next closest boundary, capped by a maximum offset value. For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches. During training, we preprocess the entire input sequence at once by constructing a time grid instead of a time cursor, allowing us to calculate boundary offsets for all tokens simultaneously. For clarity, the initial list of generated tokens is shown as empty; however, in practice, we begin with a START token.
The ON tokens that appear after a CHORD token and before the next TIMESHIFT token are considered notes of the generated chord. To make the generated chord more distinctive, we increase the velocity of these notes. Music generation continues until the time cursor reaches the duration of the video. After synthesizing the generated MIDI into an audio waveform, we apply a 3-second fade-out effect to prevent an abrupt ending and ensure the music matches the duration of the input video.
Algorithm 1 Creating boundary offsets during music generation in inference.
Input: List of input boundaries (in seconds), $\mathbf{b}$; video duration, $d$; valence, $x_{v}$; arousal $x_{a}$
Parameter: Sensitivity, $\xi$; maximum offset, $\delta_{max}$; generated boundary mask, $\mathbf{m_{b}}$; token type $t_{t}$; token value, $t_{v}$; time cursor, $c$; generation function including forward-pass and sampling, g
Output: List of generated tokens $\mathbf{t}$; list of boundary offsets, $\boldsymbol{\delta}$
 Let $c=0$; $\boldsymbol{\delta}=\left[\hskip 2.0pt\right]$; $\mathbf{t}=\left[\hskip 2.0pt\right]$.
while $c<d$ do
  # generate token as (type, value):
  $(t_{t},t_{v})=g(\mathbf{t},\boldsymbol{\delta},x_{v},x_{a})$
  if $t_{t}$ == TIMESHIFT then
   $c=c+t_{v}$
  else if $t_{t}$ == CHORD then
   $\mathbf{m_{b}}=\left|t-\mathbf{b}\right|<\xi$.
   $\mathbf{b}\left[\mathbf{m_{b}}\right]=+\infty$.
  end if
[…]
end while
3.2 Scene detection
During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations. As scene detection is a solved problem, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/. As we generate musical chords near scene boundaries, we aim to avoid an excessive number of densely spaced chords. We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than 4 s apart. The remaining timestamps serve as the input boundaries for the music generator.
3.3 Emotion mapping
We use a pretrained emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories (Sulun, Viana, and Davies 2024). Our music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories. Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane (Russell 1980) before feeding them into the music generator.
Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories (1977). They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion. We extracted and used the values corresponding to Ekman’s six basic emotions (1971), which are presented in Table 1.
Valence Arousal
Emotion Mean SD Mean SD
anger -0.51 0.20 -0.59 0.29
disgust -0.60 0.20 -0.35 0.41
fear -0.64 0.20 -0.60 0.32
joy -0.76 0.22 -0.48 0.26
sadness -0.63 0.23 -0.27 0.34
surprise -0.40 0.30 -0.67 0.27
Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions (1971), obtained from the user studies of Russell and Mehrabian (1977).
Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions. Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means. Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories. Using this parameter, we compute a single scaling coefficient applied to all valence-arousal distributions. This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
Although not used in our work, it is possible to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation. This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood.
4 Experimental setup
We perform objective and subjective evaluations of our overall model using two datasets, against the open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music (Kang, Poria, and Herremans 2024) and CMT (Di et al. 2021).
4.1 Implementation details
Our music generator consists of 11 layers, 8 attention heads, and a model dimensionality of 512. To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT (Di et al. 2021) and 33M in Video2Music (Kang, Poria, and Herremans 2024). Training is done with a context length of 1216 and a batch size of 64. The model is implemented in PyTorch (Paszke et al. 2019) and trained on a single NVIDIA A100 80GB GPU.
For video-based inference in evaluation, we set the maximum absolute value of the emotion category means to 0.8. To ensure consistency, all methods generate MIDI outputs that are synthesized into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com, followed by peak normalization to -3 dB. The audio is then merged with the input video by overwriting its audio channel. For all methods, outputs are generated in a single run across all datasets using a fixed random seed.
We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps. We use the Adam optimizer with gradient clipping to a norm of 1 (Kingma and Ba 2015). For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive. We reserve 5% of the training data for validation. Hyperparameters are selected via grid search to minimize validation loss.
4.2 Evaluation datasets
For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation (Thao, Roig, and Herremans 2023). We exclude auditory features during the emotion classification to prevent data leakage. To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music (Hussain et al. 2017). As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation. (Zander 2006).
We use the full EmoMV-C validation split, which contains 48 thirty-second videos. We filter the Ads dataset to match the total duration of EmoMV-C. We select videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad (Yang et al. 2008). We then remove videos shorter than 1 minute. To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube. We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in 24 evaluation videos. Finally, we trim the videos to a uniform duration of 1 minute. The resulting evaluation videos are then fed into the compared models.
4.3 Objective evaluation
We first evaluate music similarity using the ground-truth audio from the EmoMV-C dataset. Although the ground truth is not in MIDI, all methods generate MIDI outputs that are synthesized to audio, enabling a fair comparison in the audio domain. We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model (Won et al. 2020). This measures how closely the distribution of semantic tags in the generated music matches that of the ground-truth. We avoid using Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation (Kilgour et al. 2019), whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks (Yariv et al. 2024), using both EmoMV-C and Ads datasets. Video peaks are detected from the mean optical flow magnitude (Horn and Schunck 1981), while audio peaks correspond to audio onsets (Böck and Widmer 2013). We report AV-Align scores using windows of one video frame (33.4 ms at 30 FPS) and one second. Each window is applied to detect local peaks in both video and audio and then used to match peaks within the same temporal range for alignment.
4.4 Subjective evaluation
Using the evaluation videos from the EmoMV-C and the Ads datasets, we generate accompanying music with the three models and create survey pages, each containing two videos. For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions. We also provide output samples from eight videos online, with model names deanonymized and each model’s output appearing in the same position for clarity. [^3] https://em-sync.github.io
We enrolled 153 participants and asked them to rank the three models based on the standard criteria used in previous works (Di et al. 2021; Kang, Poria, and Herremans 2024): Music Richness (MR) — the richness and diversity of the music, independent of the video content; Music Quality (MQ) — the overall quality of the music, independent of the video content; Emotion Match (EM) — how well the music matches the video in terms of emotion; Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing; and Overall Match (OM) — how well the music matches the video as a whole. Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst).
5 Results and conclusion
Table 2 presents objective evaluation results for EMSYNC, Video2Music (Kang, Poria, and Herremans 2024), and CMT (Di et al. 2021), on the EmoMV-C and Ads datasets. EMSYNC outperforms both baselines across a majority of the metrics. On EmoMV-C, it achieves the best frame-level alignment and the lowest KL divergence, indicating superior synchronization to video and similarity to ground-truth music. While Video2Music performs best on second-level alignment, EMSYNC remains competitive. On the Ads dataset, EMSYNC obtains the best scores across both alignment metrics, further demonstrating its robustness across different video domains. We don’t calculate KL divergence because the Ads dataset doesn’t include ground truth music.
Table 3 demonstrates the subjective results in terms of average user rankings with 95% confidence margins. A lower average score represents a better ranking, with 1 being the best and 3 the worst possible. EMSYNC achieves the best rankings across all metrics and datasets. The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video. For overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations. By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.
Dataset EmoMV-C Ads
Metric AVf AVs KL AVf AVs
CMT 0.014 0.150 4.728 0.008 0.099
Video2Music 0.010 0.245 9.789 0.002 0.107
EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150
Table 2: Objective results: AVf and AVs denote audio-video alignment in frame and second-levels respectively, where higher is better. KL denotes KL-divergence where lower is better.
D M CMT Video2Music
EMSYNC
EmoMV-C
MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
MQ 2.226 ± 0.099 1.952 ± 0.116 1.822 ± 0.113
EM 2.111 ± 0.100 2.043 ± 0.115 1.846 ± 0.117
TM 2.058 ± 0.098 2.067 ± 0.115 1.875 ± 0.120
OM 2.106 ± 0.100 2.087 ± 0.114 1.808 ± 0.116
Ads
MR 1.949 ± 0.153 2.439 ± 0.155 1.612 ± 0.140
MQ 2.469 ± 0.150 1.888 ± 0.144 1.643 ± 0.153
EM 2.214 ± 0.153 2.051 ± 0.171 1.735 ± 0.154
TM 2.245 ± 0.168 2.010 ± 0.159 1.745 ± 0.150
OM 2.255 ± 0.153 2.122 ± 0.166 1.622 ± 0.145
Table 3: Subjective results: Mean ranks ± 95% confidence margins for the subjective metrics, where 1 is the best possible and 3 is the worst possible ranking. The first two columns represent the dataset and metric, indicated by the headers D and M. The table is transposed for compactness.